# KAGNNs: Kolmogorov-Arnold Networks meet Graph Learning

### v1

@@ Abstract
In recent years, Graph Neural Networks (GNNs) have become the de facto tool for learning node and graph representations. Most GNNs typically consist of a sequence of neighborhood aggregation (a.k.a., message passing) layers. Within each of these layers, the representation of each node is updated from an aggregation and transformation of its neighbours representations at the previous layer. The upper bound for the expressive power of message passing GNNs was reached through the use of MLPs as a transformation, due to their universal approximation capabilities. However, MLPs suffer from well-known limitations, which recently motivated the introduction of Kolmogorov-Arnold Networks (KANs). KANs rely on the Kolmogorov-Arnold representation theorem, rendering them a promising alternative to MLPs. In this work, we compare the performance of KANs against that of MLPs in graph learning tasks. We perform extensive experiments on node classification, graph classification and graph regression datasets. Our preliminary results indicate that while KANs are on-par with MLPs in classification tasks, they seem to have a clear advantage in the graph regression tasks.
@@ 3.2 The KAGCN Layer
GCN-based architectures have achieved great success in node classification tasks. While in our experiments we evaluate KAGIN on node classification datasets, the objective advantage of GCN over GIN on some them does not facilitate a fair estimation of KANs’ potential in this context. To this end, we also propose a variant of the GCN model. Specifically, we substitute the parameters and ReLU function of the standard GCN [20] model with a single KAN layer (defined in Equation (6)) to obtain the KAGCN layer:
@@ 4.1 Node classification
0.7447 $\pm$ 0.0638
Citeseer
0.6837
$\pm$ 0.0117
0.4748 $\pm$ 0.0574 0.6409 $\pm$ 0.0185 0.6680 $\pm$ 0.0110
Ogbn-arxiv 0.2450 $\pm$ 0.0806 0.2396 $\pm$ 0.0718 0.2997 $\pm$ 0.0274
0.5346
$\pm$ 0.0581
@@ Results.
The results are given in Table 1. We can see that KAGIN outperforms GIN on all but one datasets. On some datasets, the difference in performance between the two models is significant. For example, on Citeseer and Cora, KAGIN offers a respective absolute improvement of $20.89\%$ and $16.17\%$ in accuracy over GIN. On the other hand, KAGCN outperforms GCN on only $2$ out of the $7$ considered datasets. On some datasets (e. g., Ogbn-arxiv), GCN significantly outperforms KAGCN. With regards to the different families of models, there is no clear winner. The KAGIN and GIN models are the best performing methods on $3$ datasets, while the KAGCN and GCN models achieve the highest accuracy on the rest of the datasets. Overall, the results in the node classification experiments are mixed.
The use of KANs in the GIN architecture brings performance improvements on $6/7$ datasets, but $5/7$ experiments overlap in the confidence intervals. Such an overlap also occurs in $4$ experiments of the GCN-based models, but overall we can contend that KAN has a more positive impact on the GIN architecture than on the GCN architecture.
@@
GIN 2 16 NA NA 5336 0.035
GIN 2 64 NA NA 27368 0.041
GIN 4 16 NA NA 6936 0.037
GIN 4 64 NA NA 52200 0.056
KAGIN 2 16 1 1 31232 0.046
KAGIN 2 64 1 1 63488 0.048
KAGIN 2 16 4 1 54656 0.059
KAGIN 2 16 4 3 70272 0.146
KAGIN 4 16 1 1 38400 0.056
KAGIN 4 64 1 1 116736 0.072
KAGIN 4 64 4 1 204288 0.109
KAGIN 4 64 4 3 262656 0.298
@@
GIN 16 NA NA 3522 0.368
GIN 64 NA NA 44802 0.372
GIN 256 NA NA 670722 0.433
GIN 512 NA NA 2652162 0.742
KAGIN 16 1 1 6752 0.474
KAGIN 32 1 1 21696 0.475
KAGIN 64 1 1 76160 0.475
KAGIN 256 1 1 1091072 0.732
KAGIN 16 1 3 10048 0.658
KAGIN 32 1 3 32384 0.659
KAGIN 64 1 3 113920 0.792
KAGIN 256 1 3 1635328 2.159
KAGIN 16 4 1 11696 0.476
KAGIN 32 4 1 37728 0.478
KAGIN 64 4 1 132800 0.485
KAGIN 256 4 1 1907456 1.074
KAGIN 16 4 3 14992 0.662
KAGIN 32 4 3 48416 0.726
KAGIN 64 4 3 170560 1.005
KAGIN 256 4 3 2451712 3.092
@@ 4.3 Graph Regression
We perform grid search to select values for the different hyperparameters. For both models, we choose the number of hidden layers from $\{2,3,4\}$, and the learning rate from $\{10^{-3},10^{-4}\}$. For GIN, we choose the hidden dimension size from $\{32,64,128,256,512,1024\}$, while for KAGIN, we choose it from $\{4,8,16,32,64,128,256\}$. For KAGIN, we also select the grid size from $\{1,3,5,8,10\}$ and the spline order from $\{3,5\}$. To produce graph representations, we use the sum operator. The emerging graph representations are finally fed to an MLP (for GIN) or a KAN (for KAGIN) layer which produces the output. We train each model for $1,000$ epochs by minimizing the mean absolute error (MAE). We use the Adam optimizer for model training [19]. We also use early stopping with a patience of $20$ epochs. For ZINC-12K, we also use an embedding layer which maps node features into $100$-dimensional vectors. We choose the configuration that achieves the lowest validation error. Once the best configuration is found, we run $10$ experiments and report the average performance on the test set. For both datasets and models, we set the number of message passing layers to $4$. On QM9, we performed a joint regression of the $12$ targets.

### v2

@@ Abstract
In recent years, Graph Neural Networks (GNNs) have become the de facto tool for learning node and graph representations. Most GNNs typically consist of a sequence of neighborhood aggregation (a.k.a., message passing) layers. Within each of these layers, the representation of each node is updated from an aggregation and transformation of its neighbours representations at the previous layer. The upper bound for the expressive power of message passing GNNs was reached through the use of MLPs as a transformation, due to their universal approximation capabilities. However, MLPs suffer from well-known limitations, which recently motivated the introduction of Kolmogorov-Arnold Networks (KANs). KANs rely on the Kolmogorov-Arnold representation theorem, rendering them a promising alternative to MLPs. In this work, we compare the performance of KANs against that of MLPs in graph learning tasks. We perform extensive experiments on node classification, graph classification and graph regression datasets. Our preliminary results indicate that while KANs are on-par with MLPs in classification tasks, they seem to have a clear advantage in the graph regression tasks. Code is available at https://github.com/RomanBresson/KAGNN.
@@ 3.2 The KAGCN Layer
GCN-based architectures have achieved great success in node classification tasks. While in our experiments we evaluate KAGIN on node classification datasets, the objective advantage of GCN over GIN on some of the datasets does not facilitate a fair estimation of KANs’ potential in this context. To this end, we also propose a variant of the GCN model. Specifically, we substitute the parameters and ReLU function of the standard GCN [20] model with a single KAN layer (defined in Equation (6)) to obtain the KAGCN layer:
@@ 4.1 Node classification
0.7566 $\pm$ 0.0718
Citeseer 0.6837 $\pm$ 0.0117 0.4748 $\pm$ 0.0574 0.6409 $\pm$ 0.0185
0.6891
$\pm$ 0.0074
Ogbn-arxiv 0.2450 $\pm$ 0.0806 0.2110 $\pm$ 0.0628
0.2997
$\pm$ 0.0274
0.2777$\pm$ 0.0699
@@ Results.
The results are given in Table 1. We can see that KAGIN outperforms GIN on all but one datasets. On some datasets, the difference in performance between the two models is significant. For example, on Citeseer and Cora, KAGIN offers a respective absolute improvement of $20.89\%$ and $20.9\%$ in accuracy over GIN. On the other hand, KAGCN outperforms GCN on $3/7$ considered datasets, albeit being the best in $2/3$ homophilic networks and exhibiting very similar accuracy in $3/4$ heterophilic ones. The KAGIN and GIN models exhibit the best performance in heterophilic networks while the KAGCN and GCN models achieve the highest accuracy on the homophilic. The use of KANs in the GIN architecture brings performance improvements on $6/7$ datasets, but $5/7$ experiments overlap in the confidence intervals, an overlap that is more prevalent with GCN-based models in $6/7$. Overall we can contend that KAN has a more positive impact on the GIN architecture than on the GCN architecture. Moreover, the introduction of KAN does not alleviate the inherent disadvantage of the model e.g. GCN accuracy with heterophilic networks [41], however, it improves the models in the majority of the results i.e. $5/7$.
@@
GIN 2 16 NA NA 5,336 0.035
GIN 2 64 NA NA 27,368 0.041
GIN 4 16 NA NA 6,936 0.037
GIN 4 64 NA NA 52,200 0.056
KAGIN 2 16 1 1 31,232 0.046
KAGIN 2 64 1 1 63,488 0.048
KAGIN 2 16 4 1 54,656 0.059
KAGIN 2 16 4 3 70,272 0.146
KAGIN 4 16 1 1 38,400 0.056
KAGIN 4 64 1 1 116,736 0.072
KAGIN 4 64 4 1 204,288 0.109
KAGIN 4 64 4 3 262,656 0.298
@@
GIN 16 NA NA 3,522 0.368
GIN 64 NA NA 44,802 0.372
GIN 256 NA NA 670,722 0.433
GIN 512 NA NA 2,652,162 0.742
KAGIN 16 1 1 6,752 0.474
KAGIN 32 1 1 21,696 0.475
KAGIN 64 1 1 76,160 0.475
KAGIN 256 1 1 1,091,072 0.732
KAGIN 16 1 3 10,048 0.658
KAGIN 32 1 3 32,384 0.659
KAGIN 64 1 3 113,920 0.792
KAGIN 256 1 3 1,635,328 2.159
KAGIN 16 4 1 11,696 0.476
KAGIN 32 4 1 37,728 0.478
KAGIN 64 4 1 132,800 0.485
KAGIN 256 4 1 1,907,456 1.074
KAGIN 16 4 3 14,992 0.662
KAGIN 32 4 3 48,416 0.726
KAGIN 64 4 3 170,560 1.005
KAGIN 256 4 3 2,451,712 3.092
@@ 4.3 Graph Regression
We perform grid search to select values for the different hyperparameters. For both models, we choose the number of hidden layers from $\{2,3,4\}$, and the learning rate from $\{10^{-3},10^{-4}\}$. For GIN, we choose the hidden dimension size from $\{32,64,128,256,512,1024\}$, while for KAGIN, we choose it from $\{4,8,16,32,64,128,256\}$. For KAGIN, we also select the grid size from $\{1,3,5,8,10\}$ and the spline order from $\{3,5\}$. To produce graph representations, we use the sum operator. The emerging graph representations are finally fed to a 2-layer MLP (for GIN) or a 2-layer KAN (for KAGIN) which produces the output. We set the batch size equal to $128$ for all models. We train each model for $1,000$ epochs by minimizing the mean absolute error (MAE). We use the Adam optimizer for model training [19]. We also use early stopping with a patience of $20$ epochs. For ZINC-12K, we also use an embedding layer which maps node features into $100$-dimensional vectors. We choose the configuration that achieves the lowest validation error. Once the best configuration is found, we run $10$ experiments and report the average performance on the test set. For both datasets and models, we set the number of message passing layers to $4$. On QM9, we performed a joint regression of the $12$ targets.